# Faster Reinforcement Learning by Freezing Slow States

### v3

@@ 11       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
[…]
which represents the revenue, variable order cost, and fixed order cost. The objective is to maximize infinite horizon discounted reward, with $\gamma=0.995$. In our experiment, we set $y_{\textnormal{max}}=50$, $\mathcal{A}_{\textnormal{order}}=\{0,5,\ldots,50\}$, $d_{t+1}=\textnormal{clip}_{[0,50]}(d_{t}+\epsilon_{t+1})$, and $\{\epsilon_{t}\}$ are i.i.d. random variables that take value 0 with probability $0.8$ and ±$\pm 1$ with probability 0.1 each. [^9] We define $\textnormal{clip}_{[a,b]}(x)=\max\{\min\{x,b\},a\}$. We use $M=M_{u}=50$ for all methods, and $M_{l}=1$ for E-FSVI because the lower-level problem is deterministic.
Figure 6: Mean return (with 95% CI) versus computational cost for the inventory problem across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and 3 times per upper-level value iteration. For the other methods, we plot 3 times per value iteration. To compute the optimal value, we ran Base E-VI for 50 iterations with $M=100$ samples per backup.
@@
Figure 7: Spatial tasks on an 11$\times$11 gridworld. Start and end locations of each task are represented by the arrows. In order to receive a reward, the agent must pick up an object from the start location of the task and move it to the end location.
@@
  • If the agent reaches the start location of the current task $i_{t}$, then $o_{t}$ transitions from 00 to $1$.
@@
Figure 8: Mean return (with 95% CI) versus computational cost for the spatial gridworld problem across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and 2 times per upper-level value iteration. For the other methods, we plot 2 times per value iteration. To compute the optimal value, we ran Base E-VI for 50 iterations with $M=100$ samples per backup.
@@
where clip is applied componentwise and $\epsilon_{t+1}\in\mathbb{R}^{3}$ and each component is i.i.d., taking value 00 with 0.9 probability, ±$\pm 0.1$ with 0.05 probability each for the $\alpha_{t}$ dimension, and ±$\pm 1$ with $0.05$ probability each for the $\eta_{t}^{+},\eta_{t}^{-}$ dimensions. The reward function is $r(s_{t},p_{t})=p_{t}\,D_{t}(p_{t},p_{t}^{\textnormal{ref}}).$
@@
Figure 9: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear architectures across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and once per upper-level value iteration. For the other methods, we plot once per value iteration.
Figure 10: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear network architectures across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and once per upper-level value iteration. For the other methods, we plot once per value iteration.
@@
    Again, we see that that the $T=3$ version of Deep-FSFVI performed poorly under both architectures, but $T=6$ and $T=12$ were both adequate.
Overall, the results suggest that frozen-state methods are broadly effective and surprisingly, seem to enjoy some additional robustness even in the face of neural network architectures that are typically difficult to fit.
@@ Appendix A Empirical Versions of VI and FSVI
2 for $k^{\prime}=1,2,\ldots,k$ do
3       for $s$ in the state space $\mathcal{S}$ do
4             $U^{k^{\prime}}(s)=\max_{a}r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k^{\prime}-1}(s^{(a,j)})$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
@@
8 for $s$ in the state space $\mathcal{S}$ do
9       $\nu^{k}(s)=\operatorname*{arg\,max}_{a}r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k}(s^{(a,j)})$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
@@
11
Algorithm 5 Empirical VI for the Base Model (Base E-VI)
1
Output: Approximation of the $T$-periodic frozen-state policy $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$ and $J_{1}^{*}$.
2 for each slow state $x\in\mathcal{X}$ do
3       for $t=T-1,T-2,\ldots,1$ do
4             for each fast state $y\in\mathcal{Y}$ do
5                   $J^{*}_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}J^{*}_{t+1}(x,y^{(a,j)})$, where $y^{(a,j)}$ is sampled from the generative model, conditional on state $(x,y)$ and action $a$.
6
@@
9
@@
13for $k^{\prime}=1,2,\ldots,k$ do
14       for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
15             $V^{k^{\prime}}(x_{0},y_{0},J^{*}_{1},\tilde{\bm{\pi}}^{*})=\max_{a}\,\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\,\Bigl{[}\tilde{R}^{(a,j)}+\gamma^{T}V^{k^{\prime}-1}(x^{(a,j)}_{T},y^{(a,j)}_{T},J^{*}_{1},\tilde{\bm{\pi}}^{*})\Bigr{]}$, where $x_{T}^{(a,j)}$, $y_{T}^{(a,j)}$, and $\tilde{R}^{(a,j)}$ are sampled using a $T$-step transition through the generative model, conditioned on $x_{0}$, $y_{0}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $J^{*}_{1}$.
@@
17
18 end for
19 for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
20       $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\Bigl{[}\tilde{R}^{(a,j)}+\gamma^{T}V^{k}(x^{(a,j)}_{T},y^{(a,j)}_{T},J^{*}_{1},\tilde{\bm{\pi}}^{*})\Bigr{]}$, where $x_{T}^{(a,j)}$, $y_{T}^{(a,j)}$, and $\tilde{R}^{(a,j)}$ are sampled as above, conditioned on $x_{0}$, $y_{0}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $J^{*}_{1}$.
21 end for
22
@@ Appendix B Implementation Details for Experimental Results
We describe the rewards used in the environment. The agent receives a reward of 2.0 for picking up the object for every task. We specify the completion reward $r(i,w)$ for task $i$ in reward state $w$ as follows. For $i=1,\ldots,4$, we have $(r(i,0),r(i,1))=(80,6)$; for $i=5,6$, we have $(r(i,0),r(i,1))=(1,1)$; and for $i=7,8$, we have $(r(i,0),r(i,1))=(2,30)$. Note that the high rewards switch from the outer ring to inner ring depending on the value of $w_{t}$, which changes state from either 00 to $1$ or $1$ to 00 with probability 0.02.
@@ Appendix C Value Function Approximation with a Linear Architecture
(Note that since $J_{T}^{*}\equiv 0$, we can set $\bm{\omega}_{T}=\bm{0}$.) The approximation $\hat{V}(\bm{\beta})$ to the upper-level value function (of the frozen-state model) $V^{*}(J_{1}^{*},\tilde{\bm{\pi}}^{*})$ is given by a parameter vector $\bm{\beta}\in\mathbb{R}^{M}$, where
@@
The interpretation of this assumption is that the feature space $\{\phi(s)\,|\,s\in\mathcal{S}\}$ lies in the convex hull of the points defined by the pre-selected states: {±$\bigl{\{}\pm(\gamma^{\prime}/\gamma)\,\phi(s_{m})\bigr{\}}_{m=1}^{M}$. To reduce notional clutter, we will define $\kappa=\gamma^{\prime}/\gamma$ to be the amplification factor induced by the features.
[…]
@@
1 for $t=T-1,T-2,\ldots,1$ do
2       for each pre-selected state $s=(x,y)\in\tilde{\mathcal{S}}$ do
3             $J_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\mathbb{E}\bigl{[}\hat{J}_{t+1}(x,f_{\mathcal{Y}}(x,y,a,w),\bm{\omega}_{t+1})\bigr{]}$.
@@
6       Set remaining entries of $J_{t}$ to zero. Update parameter vector: $\bm{\omega}_{t}^{*}=\Phi^{\dagger}J_{t}$.
7
@@
9Let $\hat{\bm{\pi}}_{\bm{\omega}^{*}}$ be greedy with respect to $\hat{J}_{t}(\bm{\omega}_{t}^{*})=\Phi\bm{\omega}_{t}^{*}$, similar to (20).
10for $i=1,2,\ldots,k$ do
11       for each pre-selected state $s_{0}\in\tilde{\mathcal{S}}$ do
12             $V^{i}(s_{0})=\max_{a}\mathbb{E}\bigl{[}\tilde{R}(s,a,\hat{J}_{1}(\bm{\omega}_{1}^{*}))+\gamma^{T}\,\hat{V}(s_{T}(a,\tilde{\bm{\pi}}_{\text{avi}}),\bm{\beta}_{i-1})\bigr{]}.$
13             Set remaining entries of $V^{i}$ to zero. Update parameter vector: $\bm{\beta}_{i}=\Phi^{\dagger}\,V^{i}$.
14
@@
18for $s_{0}$ in the state space $\mathcal{S}$ do
19       $\hat{\mu}_{(\bm{\beta}^{k},\,\bm{\omega}^{*})}(s_{0})=\operatorname*{arg\,max}_{a}\mathbb{E}\bigl{[}\tilde{R}(s_{0},a,\hat{J}_{1}(\bm{\omega}_{1}^{*}))+\gamma^{T}\hat{V}(s_{T}(a,\tilde{\bm{\pi}}_{\bm{\omega}^{*}}),\bm{\beta}_{k})\bigr{]}$.
@@ Appendix D Proofs from Sections 3 and 4
Consider a value function $V:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$. Suppose there exists $L_{V}>0$ such that for any states $(x,y)$ and $(\tilde{x},\tilde{y})$, it holds that $|V(x,y)-V(\tilde{x},\tilde{y})|\leq L_{V}\|(x,y)-(\tilde{x},\tilde{y})\|_{2}$. Then,
$\bigl{|}(H^{t}$ $V)(x,y)-(\tilde{H}^{t}V)(\tilde{x},\tilde{y})\bigr{|}$
@@
Let $(x^{\prime},y^{\prime})=(f_{\mathcal{X}}(x,y,a,w),f_{\mathcal{Y}}(x,y,a,w))$ and $(\tilde{x}^{\prime},\tilde{y}^{\prime})=(f_{\mathcal{X}}(\tilde{x},\tilde{y},a,w),f_{\mathcal{Y}}(\tilde{x},\tilde{y},a,w))$ be one-step transitions starting from $(x,y)$ and $(\tilde{x},\tilde{y})$, according to the true system dynamics. For $t=1$:
$\bigl{|}(H$ $V)(x,y)-(\tilde{H}V)(\tilde{x},\tilde{y})\bigr{|}$
$=\Bigl{|}\max_{a}\Bigl{\{}r(x,y,a)+\gamma\,\mathbb{E}[V(x^{\prime},y^{\prime})]\Bigr{\}}-\max_{\tilde{a}}\Bigl{\{}r(\tilde{x},\tilde{y},\tilde{a})+\gamma\,\mathbb{E}[V(\tilde{x},\tilde{y}^{\prime})]\Bigr{\}}\Bigr{|}$
$\leq L_{r}\bigl{\|}(x,y)-(\tilde{x},\tilde{y})\bigr{\|}_{2}+\gamma\max_{a}\mathbb{E}\bigl{|}V(x^{\prime},y^{\prime})-V(\tilde{x},\tilde{y}^{\prime})\bigr{|}$
@@
which verifies the base case. Let us now assume that (23) holds for $t-1$.
$\bigl{|}(H^{t}$ $V)(x,y)-(\tilde{H}^{t}V)(\tilde{x},\tilde{y})\bigr{|}$
$=\Bigl{|}\max_{a}\Bigl{\{}r(x,y,a)+\gamma\,\mathbb{E}[(HV)^{t-1}(x^{\prime},y^{\prime})]\Bigr{\}}-\max_{\tilde{a}}\Bigl{\{}r(\tilde{x},\tilde{y},\tilde{a})+\gamma\,\mathbb{E}[(\tilde{H}V)^{t-1}(\tilde{x},\tilde{y}^{\prime})]\Bigr{\}}\Bigr{|}$
$\leq L_{r}\bigl{\|}(x,y)-(\tilde{x},\tilde{y})\bigr{\|}_{2}+\gamma\max_{a}\mathbb{E}\bigl{|}(HV)^{t-1}(x^{\prime},y^{\prime})-(\tilde{H}V)^{t-1}(\tilde{x},\tilde{y}^{\prime})\bigr{|}$

### v4

@@ A General Lemma
$r(s_{t},a_{t})=p\,\mathbb{E}\bigl(\min\{y_{t},d_{t+1}\}\bigr)-c\,a_{t}-K\,\mathbf{1}_{\{a_{t}\,>\,0\}},$
which represents the revenue, variable order cost, and fixed order cost. The objective is to maximize infinite horizon discounted reward, with $\gamma=0.995$. In our experiment, we set $y_{\textnormal{max}}=50$, $\mathcal{A}_{\textnormal{order}}=\{0,5,\ldots,50\}$, $d_{t+1}=\textnormal{clip}_{[0,50]}(d_{t}+\epsilon_{t+1})$, and $\{\epsilon_{t}\}$ are i.i.d. random variables that take value 0 with probability $0.8$ and $\pm 1$ with probability 0.1 each. [^9] We define $\textnormal{clip}_{[a,b]}(x)=\max\{\min\{x,b\},a\}$. We use $M=M_{u}=50$ for all methods, and $M_{l}=1$ for E-FSVI because the lower-level problem is deterministic.
Figure 6: Mean return (with 95% CI) versus computational cost for the inventory problem across five independent replications. For E-FSVI, we plot the performance of the policy after each lower-level value iteration and 3 times per upper-level value iteration. For the other methods, we plot 3 times per value iteration. To compute the optimal value, we ran Base E-VI for 50 iterations with $M=100$ samples per backup.
@@
Figure 7: Spatial tasks on an $\times11$ gridworld. Start and end locations of each task are represented by the arrows. In order to receive a reward, the agent must pick up an object from the start location of the task and move it to the end location.
@@
  • If the agent reaches the start location of the current task $i_{t}$, then $o_{t}$ transitions from 0 to $1$.
@@
Figure 8: Mean return (with 95% CI) versus computational cost for the spatial gridworld problem across five independent replications. For E-FSVI, we plot the performance of the policy after each lower-level value iteration and 2 times per upper-level value iteration. For the other methods, we plot 2 times per value iteration. To compute the optimal value, we ran Base E-VI for 50 iterations with $M=100$ samples per backup.
@@
where clip is applied componentwise and $\epsilon_{t+1}\in\mathbb{R}^{3}$ and each component is i.i.d., taking value 0 with 0.9 probability, $\pm 0.1$ with 0.05 probability each for the $\alpha_{t}$ dimension, and $\pm 1$ with $0.05$ probability each for the $\eta_{t}^{+},\eta_{t}^{-}$ dimensions. The reward function is $r(s_{t},p_{t})=p_{t}\,D_{t}(p_{t},p_{t}^{\textnormal{ref}}).$
@@
Figure 9: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear architectures across five independent replications. For E-FSVI, we plot the performance of the policy after each lower-level value iteration and once per upper-level value iteration. For the other methods, we plot once per value iteration.
Figure 10: Mean return (with 95% CI) versus computational cost for the dynamic pricing problem using linear network architectures across five independent replications. For E-FSVI, we plot the performance of the policy after each lower-level value iteration and once per upper-level value iteration. For the other methods, we plot once per value iteration.
@@
    Again, we see that the $T=3$ version of Deep-FSFVI performed poorly under both architectures, but $T=6$ and $T=12$ were adequate.
Overall, the results suggest that frozen-state methods are broadly effective, and surprisingly, seem to enjoy some additional robustness even in the face of neural network architectures that can be difficult to fit.
@@ Conclusion
Funding
This research is based upon work supported by the U.S. National Science Foundation under Grant No. 1807536.
@@ Appendix A Empirical Versions of VI and FSVI
2
31exfor $k^{\prime}=1,2,\ldots,k$ do
4
5  0.5em for $s$ in the state space $\mathcal{S}$ do
6
7     0.5em $U^{k^{\prime}}(s)=\max_{a}r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k^{\prime}-1}(s^{(a,j)})$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
8     0.5em
@@
10
11 end for
12
130.5em for $s$ in the state space $\mathcal{S}$ do
14
15  0.5em $\nu^{k}(s)=\operatorname*{arg\,max}_{a}r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k}(s^{(a,j)})$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
16  0.5em
17 end for
Algorithm 5 Empirical VI for the Base Model (Base E-VI)
1
[…]
2
31exfor each slow state $x\in\mathcal{X}$ do
4
5  0.5em for $t=T-1,T-2,\ldots,1$ do
@@
7     0.5em for each fast state $y\in\mathcal{Y}$ do
8
9        0.5em $J^{*}_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}J^{*}_{t+1}(x,y^{(a,j)})$, where $y^{(a,j)}$ is sampled from the generative model, conditional on state $(x,y)$ and action $a$.
10
12        0.5em
@@
14
@@
16
@@
18
191exfor $k^{\prime}=1,2,\ldots,k$ do
20
21  0.5em for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
22
[…]
24     0.5em
@@
28
290.5em for $s_{0}=(x_{0},y_{0})$ in the state space $\mathcal{X}\times\mathcal{Y}$ do
30
[…]
32  0.5em
@@
34
@@ Appendix B Implementation Details for Experimental Results
We describe the rewards used in the environment. The agent receives a reward of 2.0 for picking up the object for every task. We specify the completion reward $r(i,w)$ for task $i$ in reward state $w$ as follows. For $i=1,\ldots,4$, we have $(r(i,0),r(i,1))=(80,6)$; for $i=5,6$, we have $(r(i,0),r(i,1))=(1,1)$; and for $i=7,8$, we have $(r(i,0),r(i,1))=(2,30)$. Note that the high rewards switch from the outer ring to inner ring depending on the value of $w_{t}$, which changes state from either 0 to $1$ or $1$ to 0 with probability 0.02.
@@ Appendix C Value Function Approximation with a Linear Architecture
[…]
@@
The interpretation of this assumption is that the feature space $\{\phi(s)\,|\,s\in\mathcal{S}\}$ lies in the convex hull of the points defined by the pre-selected states: $\bigl\{\pm(\gamma^{\prime}/\gamma)\,\phi(s_{m})\bigr\}_{m=1}^{M}$. To reduce notional clutter, we will define $\kappa=\gamma^{\prime}/\gamma$ to be the amplification factor induced by the features.
Following Tsitsiklis and Van Roy (1996), we let $\Phi\in\mathbb{R}^{N\times M}$ be a matrix with the $s$-th row equal to $\bm{\phi}^{\intercal}(s)$ and let $L\in\mathbb{R}^{M\times M}$ be a matrix with the $m$-th row equal to $\bm{\phi}^{\intercal}(s_{m})$. If we let $G$ be the remaining rows of $\Phi$, then we see that $\Phi=\bigl[L;G\bigl]$. Next, by Assumption 3, the matrix $L$ has a unique matrix inverse $L^{-1}\in\mathbb{R}^{M\times M}$. We define $\Phi^{\dagger}\in\mathbb{R}^{M\times N}$ as follows: for $m\in\{1,2,\ldots,M\}$, suppose the $m$-th column of $\Phi^{\dagger}$ is equal to the $m$-th column of $L^{-1}$, and let all the other entries of $\Phi^{\dagger}$ be zero. In other words, $\Phi^{\dagger}=[L^{-1}\;0]$. Therefore, we see that $\Phi^{\dagger}$ is a left inverse of $\Phi$:
$\Phi^{\dagger}\Phi=[L^{-1}\;0]\begin{bmatrix}L\\ G\end{bmatrix}=L^{-1}L=\mathbf{I},$
@@
2
31exfor $t=T-1,T-2,\ldots,1$ do
@@
5  0.5em for each pre-selected state $s=(x,y)\in\tilde{\mathcal{S}}$ do
6
7     0.5em $J_{t}(x,y)=\max_{a}r(x,y,a)+\gamma\,\mathbb{E}\bigl[\hat{J}_{t+1}(x,f_{\mathcal{Y}}(x,y,a,w),\bm{\omega}_{t+1})\bigr]$.
8
9     0.5em
@@
11
12  0.5em Set remaining entries of $J_{t}$ to zero. Update parameter vector: $\bm{\omega}_{t}^{*}=\Phi^{\dagger}J_{t}$.
13
@@
15Let $\hat{\bm{\pi}}_{\bm{\omega}^{*}}$ be greedy with respect to $\hat{J}_{t}(\bm{\omega}_{t}^{*})=\Phi\bm{\omega}_{t}^{*}$, similar to (20).
16for $i=1,2,\ldots,k$ do
17
18  0.5em for each pre-selected state $s_{0}\in\tilde{\mathcal{S}}$ do
19
20     0.5em $V^{i}(s_{0})=\max_{a}\mathbb{E}\bigl[\tilde{R}(s,a,\hat{J}_{1}(\bm{\omega}_{1}^{*}))+\gamma^{T}\,\hat{V}(s_{T}(a,\tilde{\bm{\pi}}_{\text{avi}}),\bm{\beta}_{i-1})\bigr].$
21
22     0.5em Set remaining entries of $V^{i}$ to zero. Update parameter vector: $\bm{\beta}_{i}=\Phi^{\dagger}\,V^{i}$.
23
@@
27for $s_{0}$ in the state space $\mathcal{S}$ do
28
29  0.5em $\hat{\mu}_{(\bm{\beta}^{k},\,\bm{\omega}^{*})}(s_{0})=\operatorname*{arg\,max}_{a}\mathbb{E}\bigl[\tilde{R}(s_{0},a,\hat{J}_{1}(\bm{\omega}_{1}^{*}))+\gamma^{T}\hat{V}(s_{T}(a,\tilde{\bm{\pi}}_{\bm{\omega}^{*}}),\bm{\beta}_{k})\bigr]$.
30  0.5em
@@ Appendix D Proofs from Sections 3 and 4
[…]
@@
[…]
@@
which verifies the base case. Let us now assume that (23) holds for t1t-1.
[…]